OpenStudy (df001):

Help solving: Cos2Θ+2Cos²Θ=2 with angle formulas

OpenStudy (df001):

I need help on factoring please

OpenStudy (df001):

but I dont know if I am suppose to bring the 2 to the other side before I factor

OpenStudy (anonymous):

What sign is that?

OpenStudy (anonymous):

What kind of math is that?

OpenStudy (anonymous):

Θ

OpenStudy (anonymous):

Is that theta?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Theta = 1/6 (6 pi n-pi), n element Z

OpenStudy (anonymous):

this is calculus isn't it

OpenStudy (anonymous):

Theta = 1/6 (6 pi n+pi), n element Z

OpenStudy (anonymous):

this is trigonometry

OpenStudy (anonymous):

haha trig is even worse for me lol

OpenStudy (df001):

Can the equation be rewritten as Cos2Θ+2Cos²Θ-2=0

OpenStudy (anonymous):

2 cos(2 Theta) = 1 alternate form

OpenStudy (anonymous):

2 cos(2 Theta)+1 = 2

OpenStudy (anonymous):

-2 sin^2(Theta)+2 cos^2(Theta)+1 = 2

OpenStudy (anonymous):

alternate form

OpenStudy (df001):

wait before factoring, do I have to bring 2 to the left side of the equation?

OpenStudy (df001):

Cos2Θ+2Cos²Θ=2

OpenStudy (df001):

If I am correct, once the two is brought to the left and factored, it should be Cos²Θ(1-sin²Θ+2-2)=0?

OpenStudy (anonymous):

I used the graph

OpenStudy (df001):

I am trying to solve this algebraically lol :)

OpenStudy (triciaal):